If the de-Broglie wavelength of the fourth Bohr orbit of hydrogen atom is 4A, the circumference of the orbit will be-

Grade:

1 Answers

Arun
5 years ago
Use mvr=nh/2π

=> mvr=nh/2π
=> 2πr=nh/mv
=> 2πr=n×lambda {lambda=h/mv}

now, n=4 and wavelength=4A

=> 2πr(circumference) = 4×4
= 16A